# Online POMDP Planning with Anytime Deterministic Optimality Guarantees

### v4

@@ 7.3 Planning for optimal action
Figure 3: The graphs show the measured planning time for RB-POMCP and DB-POMCP to find the optimal action under different UCT coefficient values. Guaranteeing the optimal action made possible by using the bounds in lemma 2. All simulation runs were capped at 3,600 seconds.
@@ Acknowledgments
Appendices
9 Mathematical Analysis
@@ 9.1 Theorem 1
$=\ldots=\mathcal{R}_{\max}\sum_{\tau=1}^{t}\delta_{\tau}.$ (50)
@@ 9.2 Lemma 1
Base case $(t=T)$ At the final time step $T$, for each belief we set the value function to be equal to the reward value at that belief state, $b_{T}$ and taking the action that maximizes the immediate reward,
@@
Induction step - We will show that the hypothesis holds for time step $t-1$. By the induction hypothesis,
@@
$\leq r(b_{t-1},a_{t-1})+\sum_{z_{t}\in\mathcal{Z}}\mathbb{P}\left(z_{t}\mid H_{t}^{-}\right)\textsc{UDB}^{\pi}(b(z_{t}))$ (64)
@@
Let $b_{t}$ denote a belief state and $\pi_{t}$ a policy at time $t$. Let ¯$\bar{\mathbb{P}}(z_{t}\mid x_{t})$ be the simplified observation model which represents the likelihood of observing $z_{t}$ given $x_{t}$. Then, the following terms are equivalent,
@@ 9.4 Theorem 2
nullifying instances of the complete probability distribution, $\mathbb{P}^{\pi}(\cdot)$, is denoted as a simplified distribution, ¯$\bar{\mathbb{P}}^{\pi}(\cdot)$. We can then split and bound from above the value function, such that the simplified value function consideres only a subset of the trajectories at time $t=0$,
@@ 9.5 Optimality Guarantees
where the first inequality is due to Theorem 2, and the second inequality is true by definition. However, the claim in Lemma 2 is a recursive claim, while the bound provided in Theorem 2 only holds with respect to the root. Thus, for completeness, we also need to show that the best action can be chosen recursively, even though the bound is ‘partial‘ in different parts of the tree.
@@
$\ \sum_{\tau_{t+1}\in\mathcal{T}(H_{t+1})}\bar{\mathbb{P}}(\tau_{t+1})-\sum_{z_{t+2}\in\bar{\mathcal{Z}}(H_{t+1},\pi_{t+1})}\sum_{\tau_{t+2}\in\mathcal{T}(H_{t+2})}\bar{\mathbb{P}}(\tau_{t+2})=0,\ \forall t\in\left[0,T-2\right].$ (111)
@@
$0\displaystyle 1-\sum_{\tau_{T}}\bar{\mathbb{P}}^{\star}(\tau_{T}\mid\tau_{t},a_{t},z_{t+1},x)=$ 0 (112)
only when all non-zero future trajectories with a prefix $(\tau_{t},a_{t},z_{t+1},x)$ have been explored. Finally, we are left to show that selecting actions based on the criteria shown in (29), results in the optimal action upon convergence. Utilizing lemma 2, the proof follows similarly to the one shown in (3.1), which concludes our derivation. ∎
10 Experiments
10.1 POMDP scenarios
@@ 10.1.1 Tiger POMDP
10.1.2 Discrete Light Dark
@@ 10.1.4 Baby POMDP
10.2 Hyperparameters

### v5

@@ 7.3 Planning for optimal action
Figure 3: The graphs show the measured planning time for RB-POMCP and DB-POMCP to find the optimal action for Rock Sample under different UCT coefficient values. Guaranteeing the optimal action made possible by using the bounds in corollary 2.1. All simulation runs were capped at 3,600 seconds.
@@ 7.4 Technical Details
8 Discussion and future work
The principal advantage of the proposed algorithm is its explicit and deterministic upper and lower bounds on the value function, quantifying the maximum deviation of the current policy from optimality. These bounds support a principled early termination by choosing a user-specified tolerance $\epsilon$, guaranteeing an $\epsilon$-optimal policy and pruning of provably irrelevant action branches - capabilities that, to our knowledge, existing online POMDP planners lack. Moreover, they allow one to utilize the bounds for deterministic limits on the probability of catastrophic outcomes, an essential feature for safety-critical tasks.
However, the present implementation has few shortcomings. First, while not limited by the theoretical derivations, our implementation considers naive bounds on the value function, that is, best-or-worst possible rewards at each unvisited history or state. This results in relatively loose bounds which hinders scalability to large POMDPs. Tighter alternatives, such as history-dependent relaxations or bounds derived from the MDP-optimal value function commonly used in other POMDP solvers would alleviate this issue. We believe that this opens a new avenue of research, that focuses on general, efficient value function bounds which are easily applicable to the present deterministic-bounds algorithm (see $V_{max,d}$ in Algorithm-$\mathcal{A}$ line 1). Second, it adds implementation complexity, especially apparent compared simple algorithms like POMCP. Last, the bounds add bookkeeping overhead in both time and memory when attached to existing algorithms, although section 6.3 shows that the asymptotic complexity can be reduced to match existing algorithms, and pruning suboptimal actions may further reduce the time efficiency.
@@ Acknowledgments
Appendix A Appendices
A.1 Mathematical Analysis
@@ A.1.1 Theorem˜1
$\ldots=\mathcal{R}_{\max}\sum_{\tau=1}^{t}\delta_{\tau}.$ (53)
@@ A.1.2 Lemma˜1
Base case $(t=T)$ - At the final time step $T$, for each belief we set the value function to be equal to the reward value at that belief state, $b_{T}$ and taking the action that maximizes the immediate reward,
@@
Induction step - We will show that the hypothesis holds for time step t1t-1. By the induction hypothesis,
@@
$\leq r(b_{t-1},a_{t-1})+\sum_{z_{t}\in\mathcal{Z}}\mathbb{P}\left(z_{t}\mid H_{t}^{-}\right)\textsc{UDB}^{\pi}(b(z_{t}))\$(67)
@@
Let $b_{t}$ denote a belief state and $\pi_{t}$ a policy at time $t$. Let $\bar{\mathbb{P}}(z_{t}\mid x_{t})$ be the simplified observation model which represents the likelihood of observing $z_{t}$ given $x_{t}$. Then, the following terms are equivalent,
@@ A.1.4 Theorem˜2
nullifying instances of the complete probability distribution, $\mathbb{P}^{\pi}(\cdot)$, is denoted as a simplified distribution, $\bar{\mathbb{P}}^{\pi}(\cdot)$. We can then split and bound from above the value function, such that the simplified value function consideres only a subset of the trajectories at time $t=0$,
@@
where the first inequality is due to Theorem 2, and the second inequality is true by definition. However, the claim in Corollary 2.1 is a recursive claim, while the bound provided in Theorem 2 only holds with respect to the root. Thus, for completeness, we also need to show that the best action can be chosen recursively, even though the bound is ‘partial‘ in different parts of the tree.
@@
[…]
@@
$1-\sum_{\tau_{T}}\bar{\mathbb{P}}^{\star}(\tau_{T}\mid\tau_{t},a_{t},z_{t+1},x)=0$ (115)
only when all non-zero future trajectories with a prefix $(\tau_{t},a_{t},z_{t+1},x)$ have been explored. Finally, we are left to show that selecting actions based on the criteria shown in (29), results in the optimal action upon convergence. Utilizing corollary 2.1, the proof follows similarly to the one shown in (2.3), which concludes our derivation. ∎
A.1.7 Gap between the upper and lower bounds.
Define
$\Delta(H_{t})\;\triangleq\;\mathcal{U}^{\pi}_{0}(H_{t})-\mathcal{L}^{\pi}_{0}(H_{t}),\qquad\Delta V(H_{t})\;\triangleq\;\mathcal{V}_{\max}(H_{t})-\mathcal{V}_{\min}(H_{t}).$
Step 1: Recursion. Subtract the two expressions in (21):
$\Delta(H_{t})$ $=\sum_{\tau_{t}\in\mathcal{T}(H_{t})}\bar{\mathbb{P}}(\tau_{t})\,\Delta V(H_{t})+\!\!\!\sum_{z_{t+1}\in\bar{\mathcal{Z}}(H_{t},\pi_{t})}\!\!\!\bigl[\Delta(H_{t+1})-\!\!\!\!\sum_{\tau_{t+1}\in\mathcal{T}(H_{t+1})}\!\!\!\bar{\mathbb{P}}(\tau_{t+1})\,\Delta V(H_{t})\bigr].$ (116)
Because $\mathcal{T}(H_{t+1})$ contains only those trajectories whose next observation is inside $\bar{\mathcal{Z}}$ and whose next state is inside the current simplified state set $\bar{\mathcal{X}}(H_{t+1})$, the probability that drops out between the two sums in (116) is precisely the mass of trajectories that leave either simplified space at step $t+1$.
Step 2: Coverage-gap term. Let
$\delta(H_{t})\;\triangleq\;\sum_{\tau_{t}\in\mathcal{T}(H_{t})}\bar{\mathbb{P}}(\tau_{t})\;-\!\!\!\sum_{z_{t+1}\in\bar{\mathcal{Z}}(H_{t},\pi_{t})}\;\sum_{\tau_{t+1}\in\mathcal{T}(H_{t+1})}\bar{\mathbb{P}}(\tau_{t+1}),$
i.e. the probability that either (i) $z_{t+1}\notin\bar{\mathcal{Z}}(H_{t},\pi_{t})$ or (ii) $z_{t+1}\in\bar{\mathcal{Z}}$ but $x_{t+1}\notin\bar{\mathcal{X}}(H_{t+1})$ given $\tau_{t}$. Then (116) rewrites as
$\Delta(H_{t})\;=\;\Delta V(H_{t})\,\delta(H_{t})+\!\!\!\sum_{z_{t+1}\in\bar{\mathcal{Z}}(H_{t},\pi_{t})}\!\!\!\Delta(H_{t+1}),$
or,
$\mathcal{U}^{\pi}_{0}(H_{t})-\mathcal{L}^{\pi}_{0}(H_{t})\;=\;\Delta V(H_{t})\,\delta(H_{t})+\!\!\!\sum_{z_{t+1}\in\bar{\mathcal{Z}}(H_{t},\pi_{t})}\!\!\!\left[\mathcal{U}^{\pi}_{0}(H_{t+1})-\mathcal{L}^{\pi}_{0}(H_{t+1})\right].$
Hence the upper-lower gap at node $H_{t}$ accumulates the local value spread $\Delta V(H_{k})$ only on steps where the trajectory exits either simplified space. If all future observations and states remain in $\bar{\mathcal{Z}}$ and $\bar{\mathcal{X}}$, the indicator is zero and the gap collapses; conversely, frequent exits slow down the tightening of the bounds.
Appendix B Experiments
B.1 POMDP scenarios
@@ B.1.1 Tiger POMDP
B.1.2 Discrete Light Dark
@@ B.1.4 Baby POMDP
B.1.5 Rock Sample
In the Rock Sample problem Smith and Simmons [2004] a rover explores a $k{\times}k$ grid containing $n$ rocks whose quality (good / bad) is initially unknown; where $n=15$ and $k=3$ in the experiments section. A state is the rover cell together with an $n$-bit vector of rock qualities. The agent can move deterministically in the four cardinal directions, execute a Sample action at its current cell, or invoke one of $n$ sensing actions. The sensing action returns the observation $\mathsf{good},\mathsf{bad}$} with an accuracy level whose determined as a function of the Manhattan distance to rock $i$; all other actions yield the null observation. Each step incurs a penalty $r_{\text{step}}<0$, sensing adds $r_{\text{sensor}}<0$, sampling yields $r_{\text{good}}>0$ if the rock is good and $r_{\text{bad}}<0$ otherwise, and reaching the exit column grants $r_{\text{exit}}>0$. These dynamics create a canonical exploration-exploitation trade-off in which the rover must decide whether to spend time and energy gathering information or head directly for the exit.
B.1.6 Navigate To Goal
This environment is a $k{\times}n$ grid (default $5{\times}5$) in which a state is the agent’s coordinate pair $s=(x,y)$. Five actions are available, Stay, East, West, North, and South. An intended move succeeds with probability $0.7$; the remaining $0.3$ is shared uniformly among the other adjacent cardinal neighbors (attempts to leave the grid leave the agent in place). Rewards are +10+10 for entering a goal cell, 10-10 for stepping on the a trap cell, and 0 elsewhere.
B.2 Hyperparameters